What is the equation in point-slope form of the line with x-intercept of 3 and y-intercept of -5.
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

What is the equation in point-slope form of the line with x-intercept of 3 and y-intercept of -5.
-----
Points given are (3,0) and (0,-5)
-------------
slope = -5/-3 = 5/3
y-intercept = -5
-------------
Equation:
y -= (5/3)x-5
------------------
Cheers,
Stan H.
=============

Answer by mananth(16949)

The x intercept is 3 so the point of intersection on the x axis is(3,0)
The y intercept is -5 so the point of intersection on the y axis is (0,-5)
Find the slope of the line using these co-ordinates
x1 y1 x2 y2
3 0 0 -5

slope m = (y2-y1)/(x2-x1)
( -5 - 0 )/( 0 - 3 )
( -5 / -3 )
m= 5/3

Plug value of the slope and point (3,0) in
Y = m x + b
0.00 = 5 + b
b= 0 5
b= -5
So the equation will be
Y = 5/3 x -5
m.ananth@hotmail.ca